OpenStudy (anonymous):

The difference between two numbers is 11. When the Larger number is added to two times the smaller number, the sum is 59. What are the two numbers?

OpenStudy (anonymous):

x-y = 11 x+2y=59 x = 11 +y hence, 11+y+2y =59 3y = 48 y = 16 back to the x eqt. x=11 + 16 x=27